# Investigating Popularity Bias Amplification in Recommender Systems Employed in the Entertainment Domain

### v2

@@ 1. Introduction and Motivation
The present work summarizes our research on investigating the amplification of popularity bias in recommender systems within the entertainment domain. Analyzing datasets from three sectors, music, movies, and animes, we demonstrate that popularity bias disproportionately impacts user groups with little interest in popular content, as they receive less accurate recommendations compared to those who favor popular items (Kowald et al., 2020, 2021; Kowald and Lacic, 2022; Kowald et al., 2023). Furthermore, this work provides deeper insights into the relationship between recommendation accuracy, algorithmic calibration quality, and the amplification of popularity bias. Finally, we propose directions for future research.
@@ 2.2. Datasets and Recommender System Algorithms Employed in the Entertainment Domain
We analyze two well-known personalized recommendation algorithms used in the entertainment section, namely user-based, k-nearest-neighbor CF (UserKNN) (herlocker1999algorithmic]) and non-negative matrix factorization (NMF) (Luo et al., 2014). For the sake of reproducibility (Semmelrock et al., 2023, 2024), the implementation details of these algorithms and the complete source code to reproduce all of our research results are available via GitHub [^2] https://github.com/domkowald/FairRecSys.
@@ 3. Results and Findings
In Figure 1b, we show the correlation between the popularity of music artists and their recommendation frequency in our Last.fm dataset. We see that both algorithms tend to favor popular music artists in their recommendation lists (Kowald et al., 2020, 2021). For the sake of space, we omit the results for MovieLens and MyAnimeList, but similar results can be obtained for these domains as well (Kowald and Lacic, 2022; Kowald et al., 2023). This means that the higher the popularity of an item, the higher is also the probability that this item is recommended. Next, we investigate if this item popularity bias also negatively influences user groups with little interest into popularity (i.e., our LowPop group).
@@
Figure 1. Correlation of music artist popularity and recommendation frequency in the Last.fm dataset. Both algorithms investigated tend to favor popular music artists (Kowald et al., 2020, 2021). Similar results can be obtained for the movie and anime domains (Kowald and Lacic, 2022; Kowald et al., 2023)
Next, Table 2 summarizes our results for the three entertainment datasets across the three user groups and the two algorithms. We see that the LowPop user group always receives the statistically significant (according to a t-test with $p<0.05$) worst accuracy results (MAE). In addition, this user group also receives the most miscalibrated (MC) and popularity biased (PL) results, which helps to better understand the connection between accuracy and these two metrics (Kowald and Lacic, 2022; Kowald et al., 2023). Only in the Last.fm dataset, the other user groups reach worse PL estimates, which we attribute to the special characteristics of this dataset, namely the high number of items (see Table1 and also (Kowald et al., 2020)).
@@
4. Conclusion and Future Research Directions
In this work, we summarize our research on investigating the amplification of popularity bias in recommender systems within the entertainment sector. Analyzing datasets from music, movies, and anime, we show that recommendation frequency increases with item popularity, disadvantaging users with little interest in popular content. Additionally, we explore the interplay between recommendation accuracy, algorithmic calibration, and the amplification of popularity bias. Future research should investigate additional domains with respect to popularity bias amplification (e.g., news (Lacic et al., 2022)), and propose robust methods for mitigating popularity bias (e.g., (Klimashevskaia et al., 2022)).
Acknowledgements. This work was supported by the FFG COMET program. A full version of this article was accepted as a habilitation at Graz University of Technology in June 2024, and is available via Arxiv (Kowald, 2024).

### v3

@@ 1. Introduction and Motivation
The present work summarizes our research on investigating the amplification of popularity bias in recommender systems within the entertainment domain. Analyzing datasets from three sectors, music, movies, and animes, we demonstrate that popularity bias disproportionately impacts user groups with little interest in popular content, as they receive less accurate recommendations compared to those who favor popular items (Kowald et al., 2020b, 2021; Kowald and Lacic, 2022; Kowald et al., 2023). Furthermore, this work provides deeper insights into the relationship between recommendation accuracy, algorithmic calibration quality, and the amplification of popularity bias. Finally, we propose directions for future research.
@@ 2.2. Datasets and Recommender System Algorithms Employed in the Entertainment Domain
We analyze two well-known personalized recommendation algorithms used in the entertainment section, namely user-based, k-nearest-neighbor CF (UserKNN) (Herlocker et al., 1999) and non-negative matrix factorization (NMF) (Luo et al., 2014). For the sake of reproducibility (Semmelrock et al., 2023, 2024), the implementation details of these algorithms and the complete source code to reproduce all of our research results are available via GitHub [^2] https://github.com/domkowald/FairRecSys.
@@ 3. Results and Findings
In Figure 1b, we show the correlation between the popularity of music artists and their recommendation frequency in our Last.fm dataset. We see that both algorithms tend to favor popular music artists in their recommendation lists (Kowald et al., 2020b, 2021). For the sake of space, we omit the results for MovieLens and MyAnimeList, but similar results can be obtained for these domains as well (Kowald and Lacic, 2022; Kowald et al., 2023). This means that the higher the popularity of an item, the higher is also the probability that this item is recommended. Next, we investigate if this item popularity bias also negatively influences user groups with little interest into popularity (i.e., our LowPop group).
@@
Figure 1. Correlation of music artist popularity and recommendation frequency in the Last.fm dataset. Both algorithms investigated tend to favor popular music artists (Kowald et al., 2020b, 2021). Similar results can be obtained for the movie and anime domains (Kowald and Lacic, 2022; Kowald et al., 2023)
Next, Table 2 summarizes our results for the three entertainment datasets across the three user groups and the two algorithms. We see that the LowPop user group always receives the statistically significant (according to a t-test with $p<0.05$) worst accuracy results (MAE). In addition, this user group also receives the most miscalibrated (MC) and popularity biased (PL) results, which helps to better understand the connection between accuracy and these two metrics (Kowald and Lacic, 2022; Kowald et al., 2023).
Only in the Last.fm dataset, the other user groups reach worse PL estimates, which we attribute to the special characteristics of this dataset, namely that the music domain is an entertainment domain with repeatedly consumed items (Moscati et al., 2023; Lex et al., 2020; Kowald et al., 2020a). Thus, in contrast to movies and animes, which are, in most cases, consumed only once or a few items, items in the music domain, such as artists or tracks, are typically consumed repeatedly. The PL metric, however, only accounts for unique consumption patterns of items, and therefore this metric provides unintuitive results in the music domain. This suggest, that a weighted variant of PL is needed, in which the popularity of an item for a given user is weighted with the number of items the user has consumed this item. We aim to implement and evaluate such an extended PL metric in our future research, which we describe in the following section.
@@
4. Conclusion, Discussion, and Future Research Directions
In this work, we have summarized our research on investigating the amplification of popularity bias in recommender systems within the entertainment sector. Analyzing datasets from music, movies, and anime, we have shown that recommendation frequency increases with item popularity, disadvantaging users with little interest in popular content. Additionally, we have explored the interplay between recommendation accuracy, algorithmic calibration, and the amplification of popularity bias. We want to further discuss this interplay and its relation to algorithmic fairness in the following paragraph.
Discussion and Relation to Algorithmic Fairness
The metrics of accuracy differences, popularity lift, and miscalibration offer complementary perspectives on how recommender systems treat different user groups, and, together, they reveal important dimensions of algorithmic fairness. Accuracy differences expose disparities in the quality of recommendations received by different user groups. When users with low interest in popular content consistently receive lower accuracy scores, this suggests a systemic bias favoring mainstream preferences. This is not just a performance issue, but a fairness concern, as some users are consistently underserved.
Popularity lift further quantifies this issue by measuring the extent to which recommender systems amplify the popularity bias already present in user profiles. A high popularity lift for certain groups indicates that the system pushes them toward more popular items than they naturally prefer, potentially marginalizing niche tastes and reinforcing majority preferences. Miscalibration offers yet another lens, since it captures the mismatch between the user’s historical preferences and the genre distribution in the recommendations. A high miscalibration value implies that the algorithm fails to reflect users’ true interests, which often coincides with lower accuracy and higher popularity lift, particularly for users with niche preferences.
Taken together, these three metrics provide a more holistic understanding of fairness. While accuracy differences reveal who is affected, popularity lift and miscalibration help explain how and why certain user groups are disadvantaged. A system that exhibits large disparities across all three dimensions can be considered unfair in its personalization strategy, as it systematically underrepresents the interests of less mainstream users. Thus, improving algorithmic fairness in recommender systems requires not only maximizing overall accuracy but also minimizing popularity lift and miscalibration, especially across diverse user groups. Balancing these metrics is essential for building systems that serve users equitably, beyond the majority preference.
Future Research Directions
Future research should investigate additional domains with respect to popularity bias amplification (e.g., news (Lacic et al., 2022)), and also investigate the usefullness of the different metrics (especially popularity lift) in domains with repeat consumption patterns such as music, as already outlined in Section 3. Furthermore, we plan to work on robust methods for mitigating popularity bias, e.g., based on calibration techniques (Klimashevskaia et al., 2022), and evaluating these methods in online user studies (in addition to the offline experiments provided in this work). Finally, we would also like to investigate further aspects of trustworthiness in recommender systems, such as privacy, and study their relation to algorithmic fairness and popularity bias (Müllner et al., 2024).
Acknowledgements.
This work was supported by the FFG COMET program. A full version of this article was accepted as a habilitation at Graz University of Technology in June 2024, and is available via Arxiv (Kowald, 2024).